# Uncertainty Quantification in Retrieval Augmented Question Answering

### v1

@@ 6 Conclusions
In this work we focus on retrieval augmented QA and present an approach to answer uncertainty prediction that relies on single passage utilities. We train a small neural model on passage utility judgements collected from the target QA model. We show that our uncertainty estimator is competitive or better than existing strong error prediction methods while being light-weight. Our experiments also show that our approach is particularly good in cases of extreme answer uncertainty such as questions about non-existing entities, where bigger QA models are prone to confidently formulate an incorrect answer. As future work, we would like to explore how to extend our approach to long-form generation tasks, such as query focused-summarisation.
7 Limitations
Instruction-tuned models are known to refuse to answer questions, i.e., they produce answers such as “This information is not available in the text”. Refusing to answer is an adequate response when none of the input passages contains appropriate information. However, in many cases QA models refuse when in fact they should provide an answer (Adlakha et al., 2024; Liu et al., 2024a). Following previous work (Farquhar et al., 2024), we did not explicitly instruct the QA models to abstain and consider all cases where the answer does not match the goldstandard as incorrect. Refusing to answer is in our setting an indication of uncertainty (i.e., the QA model cannot provide a correct answer) which we aim to predict.
In this work, we focus on answer uncertainty estimation for short-form information-seeking QA tasks (Rodriguez and Boyd-Graber, 2021) where the answer can be often found in one Wikipedia passage. Going forward, it would make sense to extend our approach to multi-hop and related questions involving more complex reasoning (Yang et al., 2018; Pal et al., 2022). Although we expect Passage Utility to be effective in estimating the usefulness of individual passages, it is also possible that a more complex Passage Utility aggregation function is required (Dong et al., 2018).
Appendix A Test Time Cost of Uncertainty Estimation Methods
@@
Table 3 shows the cost of estimating uncertainty for question $x$, measured by the number of inference calls required. Simple information theoretic methods (e.g., PPL) require a single call to the target QA model with the retrieval augmented QA prompt (i.e., $|R|$ retrieved passages and question $x$). However, approaches that estimate uncertainty based on diversity (e.g., Regular Entropy, Semantic Entropy, and p(true)) require generating $N$ answers, i.e., $N$ inference calls with the retrieval augmented QA prompt. In addition, Semantic Entropy requires the computation of answer clusters (i.e., grouping answers with the same meaning), so additional calls to an entailment model are required to compare the set of sampled answers. p(true) requires one additional LLM call to elicit a True/False answer but with a very long prompt including in-context examples and the assessment question with the $|R|$ retrieved passages, sampled and most likely answers, and question $x$ (see Table 7). In contrast, our approach requires $|R|$ utility predictions with a BERT-sized model.
@@
Table 3: Number and type of inference calls required to estimate answer uncertainty for question $x$ and set of retrieved passages $R$. G means inference is performed with a retrieval augmented QA model, i.e., a LLM forward pass with the prompt including the set of $|R|$ retrieved passages and question $x$ to generate a candidate answer $y$. E is inference with an evaluation model, e.g., a forward pass to ask an LLM for correctness in p(true) or a forward pass with an entailment model in Semantic Entropy. Bert-F is an inference call to predict passage utility for passages $p$ in $R$ and question $x$.
Appendix B Experimental Details
B.1 Datasets and Splits
Dataset Train Dev Test Example Question Example Answer
PopQA 10,000 1,267 3,000 Who was the director of Champion? Rabi Kinagi
RefuNQ 2,173 Who does the voice over in the Requirtion?
@@ B.2 Implementation Details
For all question answering tasks, we use the off-the-shelf Contriever-MSMARCO (Izacard et al., 2022) tool to retrieve sets of passages $R$ for question $x$ from Wikipedia and the official Wikipedia embeddings based (2018 snapshot) as our document knowledge-base. For PopQA, we follow the work by Mallen et al. (2023) who also use the full 2018 English Wikipedia dump.
@@
We train a different predictor for each target QA model and QA task. Given the large number of predictors required in our experiments, we initially tested the hyper-parameters used in Fang et al. (2024) on the NQ dataset and choose a set thereof for all predictor instances. We train each predictor for 3 epochs, with a batch size of 32 examples, learning rate equal to 2e-5, and weight decay 0.001. For each predictor we performed search on values for $\lambda$, i.e., the contribution of the $\mathcal{L}_{BCE}$ loss (Equation 4), and different criteria for model selection, i.e., the best at pairwise ranking or at both pairwise ranking and accuracy prediction (combined).
Table 5 shows the configuration for each predictor. Table cells show selection criteria (R for ranking and C for combined) and the value for $\lambda$. A trend that seems to emerge for Llama-3.1-8B and Mistral-7B-v0.3 is that best predictors tend to rely more on the target QA model accuracy, potentially indicating that their answers in some cases depend less on context. However, for the smaller model Gemma2-2B all predictors achieve a good ranking, which supports the hypothesis that small models rely more on the provided content to formulate their answers. At inference time we predict a single Passage Utility score given by the selected best checkpoint. For all predictor instances (except for all WebQ and PopQA predictors and the predictor for Llama-3.1-8B and NQ), we use half of the available training data to speed up experiments. Training and inference was run on a single A100-40GB GPU, training takes less than 12 hours depending on the dataset.
Models NQ TQA WebQ SQuAD PopQA
Gemma2.9B R, 0.25 R, 0.25 C, 1 C, 1 R, 0.25
Llama-3.1-8B C, 0.25 C, 1 R, 0.25 C, 1 C, 1
Mistral-7B-v0.3 R, 0.25 C, 1 R, 0.25 C, 1 C, 1
Gemma2.2B R, 0.25 R, 0.25 R, 0.25 R, 0.25 R, 0.25
Gemma2.27B R, 0.25 C, 1 C, 1 R, 0.25 R, 0.25
@@
In the additional results section of the appendix (Section C.3), we report the following additional answer uncertainty estimation methods. Maximum Sequence Probability (MSP) is based on the probability of the most likely answer and is computed as
@@
Note that, in contrast to $PPL(x,R,\mathcal{M})$ reported in the main section of the paper, this metric is biased by answer length, i.e., identifying an answer to have low probability (low confidence) because of its length. Despite the fact that QA models are instructed to produce short answers, they do not always follow instructions. For this reason, we consider perplexity a more accurate metric. Indeed, the length of the answer could be a feature indicating that the model is uncertain about the answer.
We estimate answer uncertainty from the Average Answer Length (Avg.Ans.Len) as the average number of words in the sampled answers. We also report Cluster Assignment (CA) which is a variant of SE without answer probabilities where the probability of each generated meaning (i.e., a cluster) is approximated from the number of answers in the cluster. We found that in general CA estimations are very close to Semantic Entropy ones.
Another uncertainty estimation approach is the negative mean Point-wise Mutual Information (PMI; Takayama and Arase 2019) over tokens; i.e., it compares the probability of answer sequence $y$ given a prompt with question $x$ and passages $R$ w.r.t the probability given by $\mathcal{M}$ to $y$ without any context. Intuitively, the higher the point-wise mutual information, the more certain the QA model is on generating $y$ (i.e., the answer is related to or depends on $x$ and $R$). PMI is computed as
$\text{PMI}(x,R,\mathcal{M})=$ (10)
$-\,\frac{1}{|y|}\sum_{t=1}^{|y|}\text{log}\dfrac{p(y_{t}|y_{1..t-1},x,R;\mathcal{M})}{p(y_{t}|y_{1..t-1};\mathcal{M})}.$
We use the implementation provided by Farquhar et al. (2024) to compute Regular Entropy, Semantic Entropy, Cluster Assignment, and p(true).
@@
We use the implementation provided by Farquhar et al. (2024) for the AUROC, Accuracy at Xof rejection, and AURAC metrics.
We use Qwen2-72B-Instruct (Yang et al., 2024) to obtain accuracy judgments (i.e., $A$ judge, Section 4); specifically, we use the Activation-aware Weight Quantization (Lin et al., 2024) version Qwen2-72B-Instruct-AWQ. We prompt the accuracy evaluator with the prompt proposed by Sun et al. (2024), as we found it to perform well. The accuracy evaluation (AccLM) prompt is shown in Table 8. In a sample of 840 generated answers human and LLM-based judgment of correctness agreed 98% of the time (Sun et al., 2024). As a reference point, to relate to accuracy as computed in previous work, we report retrieval augmented QA accuracy (Acc) defined as whether the gold answer is contained in the generated answer (Mallen et al., 2023; Asai et al., 2024).
@@
Table 7: Prompt used for the p(true) comparison approach. The items in blue are filled in with in-context examples from the training set and the current example being evaluated. $N$ represents the number of sampled answers. The ’sequence of in-context examples’ prefix is a sequence of examples taken from the training split with the same question format but with the answer to The possible answer is: resolved.
@@
Accuracy evaluation (AccLM) prompt You need to check whether the prediction of a question-answering system to a question is correct. You should make the judgment based on a list of ground truth answers provided to you. Your response should be "correct" if the prediction is correct or "incorrect" if the prediction is wrong. Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"] Prediction: W Shakespeare Correctness: correct Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"] Prediction: Roma Gill and W Shakespeare Correctness: correct Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"]" Prediction: Roma Shakespeare Correctness: incorrect Question: What country is Maharashtra Metro Rail Corporation Limited located in? Ground truth: ["India"] Prediction: Maharashtra Correctness: incorrect Question: What’s the job of Song Kang-ho in Parasite (2019)? Ground truth: ["actor"] Prediction: He plays the role of Kim Ki-taek, the patriarch of the Kim family. Correctness: correct Question: Which era did Michael Oakeshott belong to? Ground truth: ["20th-century philosophy"] Prediction: 20th century." Correctness: correct Question: Edward Tise (known for Full Metal Jacket (1987)) is in what department? Ground truth: ["sound department"] Prediction: 2nd Infantry Division, United States Army Correctness: incorrect Question: What wine region is Finger Lakes AVA a part of? Ground truth: ["New York wine"] Prediction: Finger Lakes AVA Correctness: incorrect Question: question Ground truth: gold answer Prediction: generated answer Correctness:
Table 8: Prompt used for LLM-based accuracy evaluation (AccLM).
B.3 Prompts

### v2

@@ 6 Conclusions
In this work we address uncertainty estimation in the context of retrieval augmented QA with a method that relies on individual passage utilities. Key in our approach is the definition of utility in terms of the behaviour of the QA model and whether it is able to provide a correct answer given a retrieved passage. We train a small neural model on passage utility judgements elicted from the QA model’s responses and use utility predictions to estimate answer uncertainty. Experimental results show that our uncertainty estimator is competitive or better than existing strong methods while being light-weight. Future work could extend this approach to long-form generation tasks (Stelmakh et al., 2022; Gao et al., 2023; Min et al., 2023; Zhang et al., 2024a) where evaluating whether answer correctness is more challenging (Zhang et al., 2024b) and to multi-modal QA scenarios (Borszukovszki et al., 2025).
Appendix A Ablation of Passage Utility Predictor Training Objective
Table 3 shows AUROC results on answer uncertainty prediction with Passage Utility estimators trained with different variants of the objective in Equation (5). The first row shows the full objective (see training details in Section E.2), the second row shows a variant where the ranking objective uses only entailment annotations ($e$), and in the third row the objective is solely based on accuracy prediction ($\mathcal{L}_{BCE}$). As can be seen, there is a drop in performance when the pairwise ranking loss is not used (i.e., last line of Table 3); this component of the objective provides a smoother signal on passage utility which is empirically beneficial. However, when the pairwise ranking loss is only based on entailment, performance drops by several points, highlighting the importance of combining both to model the QA answering behaviour.
Objective Terms G9B L8B M7B
$\mathcal{L}_{rank},(e+a)/2+\lambda\,\mathcal{L}_{BCE}$ 0.80 0.81 0.80
$\mathcal{L}_{rank},(e)+\lambda\,\mathcal{L}_{BCE}$ 0.73 0.73 0.71
$\mathcal{L}_{BCE}$ 0.78 0.78 0.80
Table 3: Answer uncertainty estimation with Passage Utility predictors trained with different variants of the objective in Equation (5). AUROC values for Gemma2-9B (G9B), Llama3.1-8B (L8B), and Mistral-7B-v0.3 (M7B) are averaged over development sets.
Appendix B Test Time Cost of Uncertainty Estimation Methods
@@
Table 4 shows the cost of estimating uncertainty for question $x$, measured by the number of inference calls required. Simple information theoretic methods (e.g., PPL) require a single call to the target QA model with the retrieval augmented QA prompt (i.e., $|R|$ retrieved passages and question $x$). However, approaches that estimate uncertainty based on diversity (e.g., Regular Entropy, Semantic Entropy, and p(true)) require generating $N$ answers, i.e., $N$ inference calls with the retrieval augmented QA prompt. In addition, Semantic Entropy requires the computation of answer clusters (i.e., grouping answers with the same meaning), so additional calls to an entailment model are required to compare the set of sampled answers. p(true) requires one additional LLM call to elicit a True/False answer but with a very long prompt including in-context examples and the assessment question with the $|R|$ retrieved passages, sampled and most likely answers, and question $x$ (see Table 10). In contrast, our approach requires $|R|$ utility predictions with a BERT-sized model. For instance, in our experimental setup with $N=10$ samples and retrieval augmented QA with $|R|=5$, the Semantic Entropy approach would require 11 forward passes with the QA model prompted with 5 passages (one the greedy candidate and 10 random samples) plus 45 calls to the entailment model. Our approach requires 5 forward passes with the BERT-based model.
@@
Table 4: Number and type of inference calls required to estimate answer uncertainty for question $x$ and set of retrieved passages $R$. G means inference is performed with a retrieval augmented QA model, i.e., an LLM forward pass with the prompt including the set of $|R|$ retrieved passages and question $x$ to generate a candidate answer $y$. E is inference with an evaluation model, e.g., a forward pass to ask an LLM for correctness in p(true) or a forward pass with an entailment model in Semantic Entropy. Bert-F is an inference call to predict passage utility for passages $p$ in $R$ and question $x$.
Appendix C Generality of the Uncertainty Aggregation Strategy
To validate the intuition behind our uncertainty estimation step (Section 3.2), we compare the behaviour of the QA model when prompted with individual passages in $R$ versus when prompted with $R$ (i.e., the top-$|R|$ passages). In particular, we want to inspect the proportion of cases where our stratege of taking the maximum uitility score among the passages in $R$ does not agree with the entire set accuracy. In other words, we are interested in cases where the QA model is accurate with at least one individual passage (IndAcc\,$\uparrow$) but answers incorrectly when promtped with $R$ and vice versa. In this study, we consider two individual passage accuracy variants (related to the combined definition of Passage Utility Section 3.1). One is based on accuracy $a$ (see Section 4) and is either 0 or 1. The other one is also based on accuracy but smoothed by entailment $e$ (computed by an off-the-shelf entailment model; see Section 4) and downgrades cases where $a=1$ into $a=0$ if $e<0.5$. The latter occurs in cases where the QA model produces an answer that is accurate but not entailed by the passage.
We analyse the behaviour of four QA models (LLMs) across five datasets in our experimental setup. Table 5 shows the proportion of model disagreements across development sets. We can see that such disagreements amount to a relatively small number in both settings, i.e., when at least one individual passage in $R$ yields a correct answer (IndAcc\,$\uparrow$) but the QA model prompted with $R$ yields an incorrect answer (RAcc\,$\downarrow$) and vice versa (RAcc\,$\uparrow$ IndAcc\,$\downarrow$). The results in Table 5 confirm that our aggregation approach based on individual passages is fairly general. It approximates answer uncertainty when prompting with $|R|$ passages while avoiding the complexity of estimating uncertainty over all possible combinations of input passages in terms of number and order.
G9B G27B L8B M7B
[…]
$a$ + $e$ 0.04 0.04 0.05 0.06
RAcc\,$\uparrow$ IndAcc\,$\downarrow$ $a$ 0.01 0.01 0.01 0.01
Table 5: Average proportion of cases (five development sets) where at least one individual passage in $R$ leads a correct answer (IndAcc\,$\uparrow$) but the QA model prompted with $R$ yields an incorrect one (RAcc\,$\downarrow$) (and vice versa RAcc\,$\uparrow$ IndAcc\,$\downarrow$).
Our study is related to the issue of understanding LLM sensitivity to external evidence (Xie et al., 2024; Liu et al., 2024b), i.e., how the type of evidence (supportive, contradictory, irrelevant, or misleading), the amount, and order of presentation affect LLM predictions and interact with parametric knowledge. The Passage Utility predictor is trained to predict the error of a target QA model (LLM) on answering questions, independent of the type of passage or any memorized knowledge. Given a question-passage pair, if the LLM relies on its memorized knowledge rather than adapting to the passage and still produces the correct answer, or conversely, adapts to the passage but produces an incorrect answer, then the Passage Utility predictor should reflect this outcome by predicting a correct or incorrect answer accordingly Xie et al. (2024).
Note that the Passage Utility predictor is meant to be synchronized with the target QA model and judgments of what is (or not) a correct answer. If an answer to a question changes, the target QA model answer correctness on this question may also change, and the Passage Utility predictor should also reflect this (i.e., it should be adapted with new examples). Multiple passage interactions studied in datasets with synthetic evidence (Longpre et al., 2021; Xie et al., 2024) are observed to a lesser extent in our experiments with six datasets and external retrievers. This has also been recently pointed out in Hagström et al. (2025). Our approach could be combined with additional features to capture more complex interactions (Dong et al., 2018). Investigating and understanding the relation between QA model uncertainty and (improving) context utilization is an interesting topic on its own right (Xie et al., 2024; Longpre et al., 2021; Hagström et al., 2025) but out of scope for this paper.
Appendix D Out-of-distribution Generalization of Uncertainty Estimation
We next assess the generalization ability of our Passage Utility estimator. Following previous work on question answering and out-of-distribution (o.o.d) scenarios (Kamath et al., 2020; Zhang et al., 2021), we train a Passage Utility predictor on the SQuAD dataset and then use it to predict zero-shot (i.e., without further fine-tuning) passage utilities on all other datasets (test set). As p(true) relies on 20 in context training examples, we also evaluate its ability to generalise in out-of-distribution settings.
Table 6 shows AUROC for answer uncertainty prediction in o.o.d scenarios. As an upper bound, the upper-block of the table shows AUROC values for the in distribution (i.i.d) scenario for Passage Utility and p(true). We compare o.o.d performance w.r.t. PPL and Semantic Entropy which do not rely on training examples. Although Passage Utility performance decreases in o.o.d settings, it remains competitive in four out of five datasets. In these four cases, it is always statistically significantly different from the PPL method and comparable to p(true) and Semantic Entropy. Interestingly, p(true)’s performance also drops in all o.o.d test sets showing that relying on a fixed number of in-context learning examples is neither a robust nor scalable adaptation method. Note that we focus on zero-shot accuracy to assess bare transfer performance. However, it would make sense to train a unified Passage Utility predictor for all QA tasks following a meta-learning scheme for better adaptation (Finn et al., 2017). In addition, the model could be continually adapted via few examples (Kamath et al., 2020; Zhang et al., 2021). However, we leave this to future work.
NQ TQA WebQ PopQA RefuNQ
p(true) (i.i.d) 0.73 0.75 0.67 0.81
Passage Utility (i.i.d) 0.76 0.85 0.69 0.86
PPL 0.64   0.68 0.52   0.59 0.51
Semantic Entropy 0.70 0.73   0.58 0.73   0.59
p(true) (o.o.d) 0.67 0.63 0.63 0.72 0.62
Passage Utility (o.o.d) 0.65 0.79 0.60 0.72 0.79
Table 6: Out-of-domain performance of Passage Utility predictor for Gemma2-9B (lower block). Best values are highlighted in bold; we also mark with  next best values which are significantly different using the paired De Long test ($p<0.05$). Uncertainty predictors are trained on SQuAD and evaluated zero-shot on NQ, TQA, WebQ, PopQA, and RefuNQ test sets. The upper block reports AUROC values from our main in-distribution experiments (Table 1).
Appendix E Experimental Details
E.1 Datasets and Splits
@@
Dataset Train Dev Test Example Question Example Answer
PopQA 10,000 1,267 3,000 Who was the director of Champion? Rabi Kinagi
RefuNQ 2,173 Who does the voice over in the Requirtion?
@@ E.2 Implementation Details
For all question answering tasks, we use the off-the-shelf Contriever-MSMARCO tool (Izacard et al., 2022) to retrieve sets of passages $R$ for question $x$ from Wikipedia and the official Wikipedia embeddings based (2018 snapshot) as our document knowledge-base. For PopQA, we follow the work of Mallen et al. (2023) who also use the full 2018 English Wikipedia dump.
@@
We train a different predictor for each target QA model and QA task. Given the large number of predictors required in our experiments, we initially tested the hyper-parameters used in Fang et al. (2024) on the NQ dataset and choose a set thereof for all predictor instances. We train each predictor for 3 epochs, with a batch size of 32 examples, learning rate equal to 2e-5, and weight decay 0.001 (with the exception of Llama-3.1-8B and WebQ where we used 0.01). For each predictor we performed search on values for $\lambda$, i.e., the contribution of the $\mathcal{L}_{BCE}$ loss (Equation 5), and different criteria for model selection, i.e., the best at pairwise ranking or at both pairwise ranking and accuracy prediction (combined).
Table 8 shows the configuration for each predictor. Table cells show selection criteria (R for ranking and C for combined) and the value for $\lambda$. At inference time we predict a single Passage Utility score given by the selected best checkpoint. For all predictor instances (except for all WebQ and PopQA predictors and the predictor for Llama-3.1-8B and NQ), we use half of the available training data to speed up experiments. Training and inference was run on a single A100-40GB GPU; training ranges from 2 to 12 hours depending on the dataset.
Models NQ TQA WebQ SQuAD PopQA HotPotQA
Gemma2-9B C, 0.25 C, 0.25 R, 0.25 C, 1 C, 1 C, 0.25
Gemma2-27B C, 0.25 C, 1 C, 1 C, 0.25 R, 0.25 C, 1
Llama-3.1-8B C, 0.25 C, 0.25 C, 0.25 C, 1 C, 1 C, 0.25
Mistral-7B-v0.3 R, 0.25 C, 1 C, 0.25 C, 0.25 C, 0.25 C, 1
@@
In this section, we describe additional answer uncertainty estimation methods (for which we present supplementary results in Section F). Maximum Sequence Probability (MSP) is based on the probability of the most likely answer and is computed as:
@@
Note that, in contrast to $\text{PPL}(x,R,\mathcal{M})$ reported in the main section of the paper, this metric is biased by answer length, i.e., identifying an answer to have low probability (low confidence) because of its length. Despite the fact that QA models are instructed to produce short answers, they do not always follow instructions. For this reason, we consider perplexity a more accurate metric. Indeed, answer length could indicate that the model is uncertain about the answer. Thus, we also estimate answer uncertainty from the Average Answer Length (AvgAnsLen) as the average number of words in the sampled answers. As seen in Section F.2, Table 14, MSP and AvgAnsLen perform similarly across the board.
We also report Cluster Assignment (CA) which is a variant of SE without answer probabilities where the probability of each generated meaning (i.e., a cluster) is approximated from the number of answers in the cluster. We found that in general CA estimations are very close to Semantic Entropy ones.
Another uncertainty estimation approach is the negative mean Point-wise Mutual Information (PMI; Takayama & Arase 2019) over tokens; i.e., it compares the probability of answer sequence $y$ given a prompt with question $x$ and passages $R$ w.r.t the probability given by $\mathcal{M}$ to $y$ without any context. Intuitively, the higher the point-wise mutual information, the more certain the QA model is on generating $y$ (i.e., the answer is related to or depends on $x$ and $R$). PMI is computed as:
$\text{PMI}(x,R,\mathcal{M})=-\,\frac{1}{|y|}\sum_{t=1}^{|y|}\text{log}\dfrac{p(y_{t}|y_{1..t-1},x,R;\mathcal{M})}{p(y_{t}|y_{1..t-1};\mathcal{M})}.$ (13)
We also report Retriever Score as a baseline for Passage Utility. Instead of using the predicted Passage Utility we use the original relevance score assigned by the external retriever (i.e., Contriever MS-MARCO).
We use the implementation provided in Farquhar et al. (2024) to compute Regular Entropy, Semantic Entropy, Cluster Assignment, and p(true). Note that we do not include the supervised baseline reported in Farquhar et al. (2024) as the authors show that it underperforms simple information-theoretic metrics and in addition only works for white-box models. Note that while AvgAnsLen and Retriever Score do not strictly provide scores in the $[0,1]$ interval, the package that computes AUROC finds ranking decision thresholds. [^3] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.det_curve.html
@@
We use the implementation provided in Farquhar et al. (2024) to compute AUROC, Accuracy at Xof rejection, and AURAC metrics.
We use Qwen2-72B-Instruct (Yang et al., 2024) to obtain accuracy judgments (i.e., $A$ judge, Section 4); specifically, we use the Activation-aware Weight Quantization (Lin et al., 2024), version Qwen2-72B-Instruct-AWQ. We prompt the accuracy evaluator with the prompt proposed in Sun et al. (2024), as we found it to perform well on our datasets. The accuracy evaluation (AccLM) prompt is shown in Table 11. In a sample of 840 generated answers human and LLM-based judgment of correctness agreed 98% of the time (Sun et al., 2024).
@@
Table 10: Prompt used for p(true) approach. Items in blue are filled with in-context examples from the training set and the current example being evaluated. $N$ represents the number of sampled answers. The “sequence of in-context examples” prefix is a sequence of examples taken from the training split with the same question format but with the answer to The possible answer is: resolved.
@@
Accuracy evaluation prompt. You need to check whether the prediction of a question-answering system to a question is correct. You should make the judgment based on a list of ground truth answers provided to you. Your response should be "correct" if the prediction is correct or "incorrect" if the prediction is wrong. Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"] Prediction: W Shakespeare Correctness: correct Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"] Prediction: Roma Gill and W Shakespeare Correctness: correct Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"]" Prediction: Roma Shakespeare Correctness: incorrect Question: What country is Maharashtra Metro Rail Corporation Limited located in? Ground truth: ["India"] Prediction: Maharashtra Correctness: incorrect Question: What’s the job of Song Kang-ho in Parasite (2019)? Ground truth: ["actor"] Prediction: He plays the role of Kim Ki-taek, the patriarch of the Kim family. Correctness: correct Question: Which era did Michael Oakeshott belong to? Ground truth: ["20th-century philosophy"] Prediction: 20th century." Correctness: correct Question: Edward Tise (known for Full Metal Jacket (1987)) is in what department? Ground truth: ["sound department"] Prediction: 2nd Infantry Division, United States Army Correctness: incorrect Question: What wine region is Finger Lakes AVA a part of? Ground truth: ["New York wine"] Prediction: Finger Lakes AVA Correctness: incorrect Question: question Ground truth: gold answer Prediction: generated answer Correctness:
Table 11: Prompt used for LLM-based accuracy evaluation.
E.3 Prompts